# GoLLIE : Annotation Guidelines improve Zero-Shot Information-Extraction

### 11 Dec 2023 v3

@@ Abstract
Large Language Models (LLMs) combined with instruction tuning have made significant progress when generalizing to unseen tasks. However, they have been less successful in Information Extraction (IE), lagging behind task-specific models. Typically, IE tasks are characterized by complex annotation guidelines which describe the task and give examples to humans. Previous attempts to leverage such information have failed, even with the largest models, as they are not able to follow the guidelines out-of-the-box. In this paper we propose GoLLIE (Guideline-following Large Language Model for IE), a model able to improve zero-shot results on unseen IE tasks by virtue of being fine-tuned to comply with annotation guidelines. Comprehensive evaluation empirically demonstrates that GoLLIE is able to generalize to and follow unseen guidelines, outperforming previous attempts at zero-shot information extraction. The ablation study shows that detailed guidelines is key for good results. Code, data and models will be made publicly available.
@@ 1 Introduction
Current LLMs have been trained to follow instructions, but they fail to follow annotation guidelines out-of-the-box. For instance, Figure 1 shows results on domain specific zero-shot Named Entity Recognition. The results of gpt-3.5-turbo when prompted with guidelines (Ashok & Lipton, 2023) are low, around 20 F1 score on Music or Politics domains. Building a system that enables high-performance zero-shot information extraction, reducing the dependence on costly human annotations, remains an open challenge.
@@
○ GoLLIE (Guideline-following Large Language Model for IE), a LLM fine-tuned to learn how to attend to the guidelines on a small set of well known IE tasks. Comprehensive zero-shot evaluation empirically demonstrates that GoLLIE outperforms the SoTA (Wang et al., 2023a) in zero-shot information extraction (see Figure 1).
@@ 3 Approach
○ GoLLIE forces the model to attend to the details in the guidelines, performing robustly on schemas not seen during training. On this section we deep dive into the details of our approach, describing how the input and output was represented and the regularization techniques used to force the model to attend to the guidelines.
@@ 4.2 Language Models and Baselines
○ GoLLIE is a fine-tuned version of Code-LLaMA Rozière et al. (2023). Other backbone LLMs, such as LLaMA (Touvron et al., 2023a), LLaMA-2 Touvron et al. (2023b) or Falcon Penedo et al. (2023) were considered during the development, however, as our approach uses code to represent the input and output, Code-LLaMA model worked better on the preliminary experiments. In order to perform fair comparisons the baseline developed in this paper is based on Code-LLaMA as well. All the development of this paper was done with the 7B parameter version of Code-LLama, but, for a scaling analysis we also trained the 13B and 34B parameter models.
@@
Our main point of comparison is Instruct-UIE (Wang et al., 2023a) as it is the approach closest to our system, but does not use guidelines. Another system considered for comparison is PromptNER (Ashok & Lipton, 2023), which proposes to prompt GPT-3.5 and T5 with definitions using Chain-of-Though in order to perform few-shot NER. Different from us, they did not fine-tune the model to attend to the guidelines. For fair comparison, we only considered the zero-shot results reported in the paper. In addition, other SoTA systems are added for comparison when results from Instruct-UIE and PromptNER are not available. Given that our systems is designed for the zero-shot scenario, the supervised experiments are intended to verify that our system does not degrade its performance. We thus we selected, for the supervised scenario, those systems among SoTA that share the most comparable setting with us.
@@ 5.2 Zero-Shot evaluation
PromptNER (Ashok & Lipton, 2023) also adds some definition information into the prompt in order to perform zero-shot NER. We compare our approach with them (represented as GPT-3.5) in Figure 1. Although their approach leverages guidelines too, our approach performs significantly better on all datasets, showing that LLMs (even with 175B parameters) struggle to follow guidelines. They solve this by adding examples in the context but still far behind on a comparable setting (T5-XXL).
@@ 7 Conclusions
○ GoLLIE, a LLM specifically fine-tuned to comply with annotation guidelines that were devised for helping humans to annotate the dataset. A comprehensive zero-shot evaluation empirically demonstrate that annotation guidelines are of great value for LLMs, as GoLLIE successfully leverages them. GoLLIE achieves better zero-shot results than previous attempts at zero-shot IE which do not leverage the guidelines, or use models not finetuned for following guidelines.
@@ Appendix G Data-contamination statement
We believe data-contamination is a relevant problem that affects the NLP evaluations nowadays, becoming more prevalent with LLMs (Dodge et al., 2021; Magar & Schwartz, 2022; Sainz et al., 2023a; b). Detecting whether a dataset was inside a LLM pretrained corpora is challenging even with the pre-training data itself. In this paper, unfortunately, we do not have access to the pre-training data used to train Code-LLaMA the backbone LLM of our model. This issue is particularly worrying for us because one big source of contamination is probably GitHub and other code repositories which are also used to upload evaluation benchmarks Dodge et al. (2021). As Code-LLaMA is trained on code, there is a chance for this particular data-leakage. However, all of our comparisons were made against our baseline, which has the same backbone LLM as GoLLIE. Even if the results were impacted, the improvements of our model over the baseline would not be affected by data-contamination as both share the same pre-training. We hope that in the future more transparency will allow to perform safer evaluations.

### 21 Feb 2024 v4

@@ Abstract
Large Language Models (LLMs) combined with instruction tuning have made significant progress when generalizing to unseen tasks. However, they have been less successful in Information Extraction (IE), lagging behind task-specific models. Typically, IE tasks are characterized by complex annotation guidelines which describe the task and give examples to humans. Previous attempts to leverage such information have failed, even with the largest models, as they are not able to follow the guidelines out-of-the-box. In this paper we propose GoLLIE (Guideline-following Large Language Model for IE), a model able to improve zero-shot results on unseen IE tasks by virtue of being fine-tuned to comply with annotation guidelines. Comprehensive evaluation empirically demonstrates that GoLLIE is able to generalize to and follow unseen guidelines, outperforming previous attempts at zero-shot information extraction. The ablation study shows that detailed guidelines is key for good results. Code, data and models are publicly available: https://github.com/hitz-zentroa/GoLLIE.
@@ 1 Introduction
Current LLMs have been trained to follow instructions, but they fail to follow annotation guidelines out-of-the-box. For instance, Figure 1 shows results on domain specific zero-shot Named Entity Recognition. The results of gpt-3.5-turbo when prompted with guidelines (Zhang et al., 2023a) are low, around 20 F1 score on Music or Politics domains. Building a system that enables high-performance zero-shot information extraction, reducing the dependence on costly human annotations, remains an open challenge.
@@ 4.2 Language Models and Baselines
Our main point of comparison is Instruct-UIE (Wang et al., 2023a) as it is the approach closest to our system, but does not use guidelines. Another system considered for comparison is PromptNER (Zhang et al., 2023a), which proposes to prompt GPT-3.5 and T5 with definitions using Chain-of-Though in order to perform few-shot NER. Different from us, they did not fine-tune the model to attend to the guidelines. For fair comparison, we only considered the zero-shot results reported in the paper. In addition, other SoTA systems are added for comparison when results from Instruct-UIE and PromptNER are not available. Given that our systems is designed for the zero-shot scenario, the supervised experiments are intended to verify that our system does not degrade its performance. We thus we selected, for the supervised scenario, those systems among SoTA that share the most comparable setting with us.
@@ 5.2 Zero-Shot evaluation
PromptNER (Zhang et al., 2023a) also adds some definition information into the prompt in order to perform zero-shot NER. We compare our approach with them (represented as GPT-3.5) in Figure 1. Although their approach leverages guidelines too, our approach performs significantly better on all datasets, showing that LLMs (even with 175B parameters) struggle to follow guidelines. They solve this by adding examples in the context but still far behind on a comparable setting (T5-XXL).
Seen vs unseen labels:
@@ 7 Conclusions
Acknowledgments
This work has been partially supported the Basque Government (Research group funding IT-1805-22). We are also thankful to several MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and ERDF A way of making Europe; (ii) Disargue (TED2021-130810B-C21) and European Union NextGenerationEU/PRTR; (iii) AWARE (TED2021-131617B-I00) and European Union NextGenerationEU/PRTR. This work has also been partially funded by the LUMINOUS project (HORIZON-CL4-2023-HUMAN-01-21-101135724). Oscar Sainz is supported by a doctoral grant from the Basque Government (PRE_2023_2_0137). Rodrigo Agerri currently holds the RYC-2017-23647 fellowship (MCIN/AEI/10.13039/501100011033 and by ESF Investing in your future).
@@ D.2 Dataset Detailts
Regarding the splits, we use the standard train, dev test splits for every dataset. In case of ACE, we follow the split provided by Lin et al. (2020). In the case of CASIE we took the first 200 instances as validation and the last 2000 as test.
@@ Appendix G Data-contamination statement
We believe data-contamination is a relevant problem that affects the NLP evaluations nowadays, becoming more prevalent with LLMs (Dodge et al., 2021; Magar & Schwartz, 2022; Sainz et al., 2023b; a). Detecting whether a dataset was inside a LLM pretrained corpora is challenging even with the pre-training data itself. In this paper, unfortunately, we do not have access to the pre-training data used to train Code-LLaMA the backbone LLM of our model. This issue is particularly worrying for us because one big source of contamination is probably GitHub and other code repositories which are also used to upload evaluation benchmarks Dodge et al. (2021). As Code-LLaMA is trained on code, there is a chance for this particular data-leakage. However, all of our comparisons were made against our baseline, which has the same backbone LLM as GoLLIE. Even if the results were impacted, the improvements of our model over the baseline would not be affected by data-contamination as both share the same pre-training. We hope that in the future more transparency will allow to perform safer evaluations.